Given utility function U= where PX = 12 Birr, Birr, PY = 4 Birr and the income of the consumer is, M= 240 Birr. A. Find the utility maximizing combinations of X and Y. B. Calculate marginal rate of substitution of X for Y (MRSX,Y) at equilibrium and interpret your result.
1
Expert's answer
2020-02-20T09:00:25-0500

"MUx\u200b=0.5\\frac{x0.5}{y0.5\u200b}"


"MU_y=0.5 \frac {x^{0.5}}{y^{0.5}}"

"{\frac{MU_x}{p_x}}=\frac {MU_y}{p_y}"

"x \times p_x+ y \times p_y=M"

y= 3x

12* x+4*y=240

x=10, y=30

D)

"\frac {\partial U}{\partial x \partial y} =\frac {0.25} {(xy)^{0.5}}"

"MRS x.y = \frac {\partial U} {\partial x \partial y} = \frac {0.25}{(10 \times 30)^{0.5}}=0.015"